OpenStudy (anonymous):

4. A ball is thrown into the air with an initial upward velocity of 60 ft/s. Its height (h) in feet after t seconds is given by the function h = –16t² + 60t + 6. What will the height be at t = 3 seconds? 35 feet 40 feet 42 feet 45 feet

OpenStudy (anonymous):

substitute the t in the equation for a 3, and work from there

OpenStudy (anonymous):

so whats the answer

OpenStudy (anonymous):

h(t) = -16t^2 + 60t + 6 At t=3 we find h(3): h(3) = -16(3)^2 + 60(3) + 6 = -16(9) + 180 + 6 = 42 ft C.

OpenStudy (anonymous):

thanks